# Zero-Shot Position Debiasing for Large Language Models

### v1

@@ 5 Results
Table 4: Overall performance (%) on language compression tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method Summarization (%)
CNN/DM Newsroom
Biased Non-Biased Biased Non-Biased
T5 22.3 11.6 35.1 20.0
RP 23.9 15.1 47.5 22.0
GenX 17.6 13.7 29.2 19.0
FT 26.9 16.8 51.1 19.8
ZOE 27.1 17.3 50.9 21.3
The overall performances of all methods on language understanding tasks, language creation tasks and language compression tasks are listed in Table 2, Table 3 and Table 4. We have three main observations from the results.
First, LLMs are susceptible to the bias in the dataset after fine-tuning. As we can see in Table 2, after fine-tuning, FT achieves 34.7% improvement on biased samples of CoQAR, but 8.6% improvement on non-biased samples. This is because LLMs can easily overfit the shortcut of the training dataset in fine-tuning, just like existing neural networks (Ko et al., 2020).
Second, ZOE can mitigate position bias significantly on almost all datasets for three types of tasks. As shown in Table 2, Table 3 and Table 4, ZOE improves the performance on non-biased samples, ranging from 1% to 2% on almost all tasks, compared to T5 and all baselines. The reason is that ZOE can leverage pre-trained T5 with lower position bias as a slave objective by MSA and multi-objective optimization.
Third, ZOE only sacrifices a small performance on biased samples when mitigating position bias. As shown in Table 4, RP achieves comparable performance to ZOE on non-biased samples of Newsroom. However, the ROUGE-L of RP drops 3.6% compared to FT on biased samples, while that of ZOE only drops 0.2%. This is because the perturbation in RP impairs the overall data quality for fine-tuning, while ZOE only prune partial objective for optimization in §3.3.
Note that some baselines achieve poor performance, sometimes even lower than T5. First, in Table 4, GenX performs worse even than T5 on biased samples of Newsroom. The reason is that the summarization datasets are abstractive and suitable for generative models, e.g., T5, while GenX (Varab and Xu, 2023) is an extractive baseline. Second, in Table 3, the performance of SG-CQG is worse than that of FT on biased samples of CoQAR. This is because SG-CQG is a three-stage question generation model focusing on improving question diversity. The selected answer span for question generation is randomly chosen from massive generated candidates.
6 Analysis
6.1 Analysis of Unsupervised Responses
[…]
Method CoQAR CANARD
Biased N-Biased Biased N-Biased
ZOE 26.7 18.8 25.9 22.4
- w/o MSA 25.8 18.0 26.1 21.5
- w/ T5-base 26.3 18.3 26.1 22.0
- w/ T5-xlarge 26.6 17.9 25.7 22.0
To verify the effect of the unsupervised responses, we conduct analyses of unsupervised responses in various sources with different qualities in Table 5. ZOE w/o MSA, ZOE w/ T5-base and ZOE w/ T5-xlarge denote ZOE using unsupervised responses without alignment, responses from T5-base and T5-xlarge, respectively. We have two observations.
First, lower response quality leads to worse performance. As we can see, with MSA to improve the quality of unsupervised responses, ZOE outperforms ZOE w/o MSA on non-biased samples. Poor-quality responses undermine the model comprehension of the task, thus leading to worse performance.
Second, using various sources of unsupervised responses may degrade model performance. ZOE w/ T5-base and ZOE w/ T5-xlarge leverage unsupervised responses from T5-base and T5-xlarge, while they both perform worse than ZOE on CANARD. We infer that responses from T5-base or T5-xlarge use different knowledge/parameters to generate questions, which mismatch with that of T5. This difference amplifies the divergence between master objective and slave objective in §3.2. Even so, ZOE still outperforms the fine-tuned T5.
6.2 Analysis of Objective Weighting
[CQA_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[Newsroom]Refer to caption [Doc2dial]Refer to caption
Figure 3: Performance (%) over each $\alpha$ of CQA on CANARD, CQG on CoQAR, summarization on Newsroom, KGC on Doc2dial. The x-axis denotes the value of $\alpha$ and the y-axis denotes the ROUGE-L on non-biased samples.
To analyze the effect of weighting on slave objective, we present the performance of ZOE using different $\alpha$ in Fig. 3. FT denotes the fine-tuned LLM. We have two observations from the results.
First, the performance of ZOE drops with the increase of the weight of unsupervised responses in multi-objective optimization. In CQA on CANARD, the ROUGE-L score of ZOE drops from 21.9% to 21.5% when ranging $\alpha$ from 0.1 to 0.5. This is because the performance on non-biased samples depends on not only the degree of bias but also the quality on training samples. Increasing $\alpha$ will reduce the position bias of all responses, yet it will hurt the quality concurrently.
Second, our proposed ZOE outperforms fine-tuned T5 under various $\alpha$. As shown in Fig. 3, the CQA performances of ZOE on CANARD all exceed 21.5% using various $\alpha$, while FT only achieves ROUGE-L value 20.8%. This demonstrates the effectiveness and robustness of ZOE in mitigating position bias.
6.3 Analysis on Training Samples
[CQA_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[Newsroom]Refer to caption [Doc2dial]Refer to caption
Figure 4: Performance (%) over different numbers of training samples of CQA on CANARD, CQG on CoQAR, summarization on Newsroom, KGC on Doc2dial. The x-axis denotes the number of training samples and the y-axis denotes the ROUGE-L on non-biased samples dataset.
We also analyze the effect of the number of training samples to verify the effectiveness of ZOE under various low-resource settings. We plot the results in Fig. 4 and have two observations from the results.
First, all methods always perform better when increasing training samples. As we can see in Fig. 4, the ROUGE-L score of FT increases from 16.0% to 18.5% on CoQAR when the number of training samples increases from 50 to 1,000. The increase of training samples can improve the generalization ability of LLMs, thus leading to better performance on non-biased samples.
Second, our proposed ZOE outperforms fine-tuned T5 under various settings on training samples. As shown in Fig. 4, the ROUGE-L scores of ZOE depicted by the orange charts are consistently higher than that of FT in blue. This is because there are always around 40% unsupervised responses after aligning for fine-tuning when ranging the number of training samples, which are enough for effectively debiasing.
6.4 Case Study
Table 6: An example of CQA on CANARD. U1, U4, and U5 are the 1st, 4th, 5th utterances in the document and T1 is the first turn utterance of the dialogue. ‘Target’ is the target answer of the document and ‘FT’ and ‘ZOE’ are the generated answers. ‘Position’ is the position of the utterance containing the answer in the document and ‘Relative position’ is the distance of grounded utterances between the current answer and the last answer.
ID
Document
U1
Gaston had worked with players at an individual level as a hitting instructor …
U4
In the six games the Blue Jays played in those places during World Series play
U5
… and Gaston was the first ever African-American manager to win a World Series.
ID
Context
Position
T1
Question: What happens in the series?
U4
T2
Question: What else did Cito do?
Model
Question
Relative position
Target
Gaston had worked with players at an individual level
-3 (U1)
FT
Gaston was the first ever African-American manager to win a World Series.
1 (U5)
ZOE
Gaston had worked with players at an individual level
-3 (U1)
Table 7: An example of CQG on CANARD. U1 - U5 are the utterances in the document and T1 is the first turn utterance of the dialogue. ‘Target’ is the target question of the document, ‘T5-1’ and ‘T5-2’ are the questions generated by pre-trained T5. ‘Position’ is the position of the utterance containing the answer to the question in the document and ‘Relative position’ is the distance of grounded utterances between the current question and the last question.
ID
Document
U1
Gautam Gambhir, born 14 October 1981, is an Indian cricketer, …
U2
Gambhir was picked up by the Delhi Daredevils franchise in the first player auction of the Indian Premier League for a price of US\$725,000 a year.
U3
He became the second highest run-scorer of the inaugural season with …
U4
He was promoted to the post of Captain of the Delhi Daredevils for IPL Season 2010.
U5
At the end of the tournament he became the only player from Delhi Daredevils to score more than 1000 runs in the IPL.
U6
In the 2011 IPL player auction, …
ID
Context
Position
T1
Question: Who was Gautam Gambhir in Indian premier league?
U4
Source
Question
Relative position
Target
What did Gautam Gambhir do as captain?
1 (U5)
T5-1
Where was Gautam Gambhir born?
-3 (U1)
T5-2
Was Gautam Gambhir in Indian premier league?
-2 (U2)
To investigate the reason for ZOE’s effectiveness, we present an example of generated responses in CQA in Table 6 and an example of unsupervised responses in CQG in Table 7.
In CQA example, ZOE can generate an answer with lower bias than FT. As shown in Table 6, FT generates an answer from U5, adjacent to U4, the utterance containing the last turn answer. In contrast, ZOE generates an answer from U1, which is far from U4. After fine-tuning with unsupervised responses from various positions, ZOE cannot easily fall into the trap of finding answers from neighboring utterances around the utterance containing the last turn answer.
In CQG example in Table 7, ZOE uses unsupervised responses which have lower bias than the target response. As we can see, target question is based on the utterance with relative position as 1, while unsupervised questions generated from T5 are based on utterances varying in the document. Fine-tuning with T5-generated unsupervised responses encourages the model to generate questions based on utterances in various document positions, thus mitigating position bias.
7 Related Work
7.1 Position Bias and Debiasing
Position bias has been explored mainly in three extractive tasks, i.e., NLI, summarization and CQA. Besides, there are also other explored extractive tasks but with relatively less attention, e.g., sentiment analysis (Ma et al., 2021). In, NLI, Gururangan et al. (2018); Poliak et al. (2018) show that the class labels are highly correlated to certain words in the hypothesis. McCoy et al. (2019) report that models always rely on word overlap between hypothesis and premise for prediction. Karimi Mahabadi et al. (2020); Du et al. (2021) train a robust NLI model in an ensemble manner by PoE. They train a hypothesis model to learn the lexical bias, which guides the NLI model to focus on other patterns in the dataset that generalize better.
There are similar findings on position bias in CQA tasks. Weissenborn et al. (2017); Sugawara et al. (2018) demonstrate that only using partial inputs is sufficient to correctly extract the answer span for the question in most cases. Ko et al. (2020) address that the absolute position of answer spans can work as a spurious bias as models trained on biased datasets fail to generalize over unseen answer positions. Shinoda et al. (2022) report that the relative position of answer spans is another clue for position bias. To mitigate position bias, Ko et al. (2020); Shinoda et al. (2022) build bias ensemble models by PoE similar to (Karimi Mahabadi et al., 2020), where they design biased models based on position-only features to guide the CQA models to leverage more semantic features for answering.
Kedzie et al. (2018); Grenander et al. (2019) find 58% of selected summary utterances come directly from the lead utterances, and models trained on these articles perform considerably worse when utterances in the article are randomly shuffled. To mitigate lead bias, a simple but effective method is to randomly shuffle the document for summarization training (Grenander et al., 2019). Then, Xing et al. (2021) uses adversarial training for debiasing. They design a position prediction module and optimize the reverse loss for position prediction, forcing the encoder to leverage non-position features. Besides, Zhu et al. (2021) leverage lead bias for unsupervised training, i.e., treating the leading utterances from a news article as the target summary and the rest as the article.
However, existing works on mitigating position bias always focus on extractive tasks. They can hardly be transferred to generative tasks as the label space in generative tasks is too large for bias estimation or adversarial training. In this paper, we focus on mitigating position bias for generative LLMs.
7.2 Bias in LLMs and Debiasing
In pre-processing, Zmigrod et al. (2019) adopt a counterfactual data augmentation (CDA) algorithm to mitigate social bias by swapping bias attribute words (e.g., he/she) in a dataset. Choi et al. (2022) modify CDA by masking the terms casual to label to force the model to learn label-invariant features.
In in-processing, Guo et al. (2022); Li et al. (2023); Yang et al. (2023) propose a two-stage adversarial method for debiasing, where they first train a continuous prompt to enlarge the bias of utterance pairs and then force the LLMs to minimize the difference of utterance pairs using the prompt.
In post-processing, Schick et al. (2021) propose a decoding algorithm that reduces the probability of a model producing biased text using a textual description of the undesired behavior for prompting. Meade et al. (2023) propose an ICL strategy which leverages non-biased demonstrations to guide the generation on safety.
However, existing methods for mitigating social bias either require the external bias knowledge for data augmentation and training adversarial prompts, or require a non-biased dataset for building demonstrations, which is unpractical in application. Differently, we propose a zero-shot framework for LLMs on mitigating position bias, without knowing any external bias knowledge or non-biased samples, which is general, simple but effective.
8 Conclusion
In this paper, we have proposed a zero-shot debiasing ZOE framework for LLMs. We adopt multi-objective optimization to mitigate position bias for LLMs, with a slave objective based on unsupervised responses. These responses are aligned by a proposed MSA module based on the target responses. Extensive experiments on five tasks and eight benchmark datasets show that ZOE outperforms existing baselines on non-biased samples while retaining performance or sacrificing little performance on biased samples. It proves that leveraging aligned unsupervised responses is a practicable solution to mitigate position bias for generative LLMs.
9 Limitations
This work has the following limitations. First, ZOE needs a pre-trained LLM to generate unsupervised responses with low bias, where these responses are still biased. Second, the final performance of ZOE depends on the quality of unsupervised responses, which are still noisy after the proposed MSA. In future work, we plan to address these issues by investigating non-biased models from other domains and model-based strategies to align unsupervised responses.
Ethical Considerations
We realize that there are risks in developing dialogue systems, so it is necessary to pay attention to the ethical issues of dialogue systems. We have used the public pre-trained LLMs, e.g., FlanT5-large, and public datasets, e.g., CANARD and CoQAR, to conduct the experiments. All models and datasets are carefully processed by publishers to ensure that there are no ethical problems.
Appendix A Analysis
Table 8: Overall performance (%) on language understanding tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method NLI (%) CQA (%)
SNLI QNLI CoQAR CANARD
ZOE 91.8 89.0 93.3 91.3 66.3 53.7 65.7 21.9
- w/o MSA 91.7 88.4 93.7 90.1 66.3 53.7 65.7 21.9
- w/ T5-base 91.7 87.7 93.1 91.0 64.3 52.3 61.6 19.9
- w/ T5-xlarge 92.2 88.9 93.9 91.2 65.1 53.6 67.1 22.4
Table 9: Overall performance (%) on language creation tasks. Boldface indicates the best results in terms of the corresponding metrics.
Method CQG (%) KGC (%)
CoQAR CANARD Doc2dial Mutual
ZOE 26.7 18.8 25.9 22.4 40.6 38.3 57.4 56.3
- w/o MSA 25.8 18.0 26.1 21.5 32.6 34.1 55.9 53.5
- w/ T5-base 26.3 18.3 26.1 22.0 42.8 36.3 60.6 55.4
- w/ T5-xlarge 26.6 17.9 25.7 22.0 44.1 39.1 58.7 56.9
Table 10: Overall performance (%) on language compression tasks. ‘N-Biased’ denotes performance on non-biased samples. Boldface indicates the best results in terms of the corresponding metrics.
@@ Appendix A Analysis
A.1 Analyses of Unsupervised Responses
Table 8, Table 9 and Table 10 demonstrate ZOE performance with different qualities of unsupervised responses. As we can see, lower response quality leads to worse performance. Besides, using various sources of unsupervised responses may improve or degrade model performance depending on the task type.
A.2 Analyses of Objective Weighting
Fig. 5 provides the overall performance on different $\alpha$. In most cases, ZOE outperforms fine-tuned T5.
A.3 Analyses on Training Samples
Fig. 6 provides the overall performance on different numbers of training samples. In most cases, ZOE outperforms fine-tuned T5.
[CQG_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[CQA_CANARD]Refer to caption [CQA_CoQAR]Refer to caption
@@ A.3 Analyses on Training Samples
Figure 5: Performance over each $\alpha$ of CQG on CANARD, CQG on CoQAR, CQA on CANARD, CQA on CoQAR, summarization on Newsroom, summarization on CNN/DM, KGC on Doc2dial, KGC on Mutual, NLI on SNLI, NLI on QNLI. The x-axis denotes the value of $\alpha$ and the y-axis denotes the ROUGE-L on non-biased samples.
[CQG_CANARD]Refer to caption [CQG_CoQAR]Refer to caption
[CQA_CANARD]Refer to caption [CQA_CoQAR]Refer to caption
@@
Figure 6: Performance over different numbers of training samples of CQG on CANARD, CQG on CoQAR, CQA on CANARD, CQA on CoQAR, summarization on Newsroom, summarization on CNN/DM, KGC on Doc2dial, KGC on Mutual, NLI on SNLI, NLI on QNLI. The x-axis denotes the number of training samples and the y-axis denotes the ROUGE-L on non-biased samples dataset.

### v3

@@ Appendix C Overall Analysis
Table 10: SOD Performance (%) using various unsupervised responses on language compression tasks. ‘N-Biased’ denotes performance on the non-biased datasets. Boldface indicates the best results in terms of the corresponding dataset.
@@
FT 26.9 16.8 51.1 19.8
C.1 Analyses of Unsupervised Responses
Table 8-10 demonstrate SOD performance using different qualities of unsupervised responses. As we can see, lower response quality leads to worse performance. Besides, using various sources of unsupervised responses always degrades model performance on non-biased datasets.
C.2 Analyses of Objective Weighting
Fig. 5 provides the overall performance over different $\alpha$. In most cases, SOD outperforms the fine-tuned LLM, i.e., FT.
C.3 Analyses on Training Samples
Fig. 6 provides the overall performance on different numbers of training samples. In most cases, SOD outperforms the fine-tuned LLM, i.e., FT.
C.4 Cases
We also present cases for CQA, CQG, summarization and KGC tasks in Table 1117. As we can see, FT always finds relevant information from utterances near the grounded utterances of the last turn utterance for CQA and CQG tasks or from the lead utterance for summarization and KGC tasks. While SOD can find relevant knowledge from any utterances in the document. Note that on the Mutual dataset, where each sample has four candidate responses as the document, FT always fails to select the relevant utterance from the last two responses.
[CANARD (CQG)]Refer to caption [CoQAR (CQG)]Refer to caption
[CANARD (CQA)]Refer to caption [CoQAR (CQA)]Refer to caption
@@ C.4 Cases
Figure 5: Performance over each $\alpha$ on all datasets. The x-axis denotes the value of $\alpha$ and the y-axis denotes the ROUGE-L score on non-biased datasets.
[CANARD (CQG)]Refer to caption [CoQAR (CQG)]Refer to caption
[CANARD (CQA)]Refer to caption [CoQAR (CQA)]Refer to caption
@@
Figure 6: Performance over different numbers of training samples on all datasets. The x-axis denotes the number of training samples and the y-axis denotes the ROUGE-L score on non-biased datasets.
Table 11: An example of CQA on CoQAR. U1–U4 are the first four utterances in the document, T1 and T2 are the 1st and 2nd turn utterances of the dialogue. ‘Target’ is the target answer of the example, ‘FT’ and ‘SOD’ are the generated answers from FT and SOD. ‘Position’ is the position of the grounded utterance of the answer in the document and ‘Relative position’ is the distance of grounded utterances between the current answer and the last turn answer.
ID Document
U1 This final war was to give thousands of colonists, including …, military experience
U2 By far the largest military action in which the United States engaged during this era was the War of 1812.
U3 With Britain locked in a major war with Napoleon’s France, its policy was to block American shipments to France
U4 The United States sought to remain neutral while pursuing overseas trade.
ID Context Position
T1 Question: What did this give the colonists?
Answer: Military experience.
U1
T2 Question: Whose side was the US on at first in the war of 1812?
Model Answer Relative position
Target At first neutral 3 (U4)
FT Britain. 2 (U3)
SOD Neutral 3 (U4)
Table 12: An example of CQG on CoQA. U1–U5 are the first five utterances in the document and T5 is the 5th turn utterance of the dialogue. ‘Target’ is the target question of the example and ‘FT’ and ‘SOD’ are the generated questions from FT and SOD. ‘Position’ is the position of the grounded utterance of the question in the document and ‘Relative position’ is the distance of grounded utterances between the current question and the last turn question.
ID Document
U1 John’s Metropolitan Area is the second largest Census Metropolitan Area (CMA) in Atlantic Canada
U3 Its name has been attributed to the feast day of John the Baptist, when John Cabot was believed to have sailed into the harbor in 1497.
U4 St. John’s is one of the oldest settlements in North America, with year-round settlement beginning sometime after 1630 and seasonal habitation long before that.
U5 It is not, however, the oldest surviving English settlement in North America or Canada …
ID Context Position
T5 Question: In what continent is it located?
Answer: North America.
U4
Model Question Relative position
Target What year did John Cabot arrive in the harbor? -1 (U3)
FT When was St. John’s founded?  0 (U4)
SOD What is the second largest CMA in Atlantic Canada? -3 (U1)
Table 13: An example of CQG on CANARD. U1–U5 are the first five utterances in the document and T1 is the 1st turn utterance of the dialogue. ‘Target’ is the target question of the example and ‘FT’ and ‘SOD’ are the generated questions from FT and SOD. ‘Position’ is the position of the grounded utterance of the question in the document and ‘Relative position’ is the distance of grounded utterances between the current question and the last turn question.
ID Document
U1 In May 2009, production company CinemaNX announced that it would distribute Me and Orson Welles itself …
U2 It opened the New Orleans Film Festival on October 9, 2009; and it was screened at the St. Louis International Film Festival in November 2009.
U3 The film was released in the US on November 25, 2009, and in the UK on December 4, 2009.
U4 IndieWIRE reported, “The do-it-yourself release of Richard Linklater’s Me and Orson Welles bluegot off to a very nice start, averaging \$15,910 from its four theaters, the highest PTA of all debuting films.”
U5 While Orson Welles is one the first examples of such a high-profile film going to the DIY route, if it proves successful, it’s going to be done a lot more in the future.
ID Context Position
T1 Question: What month was Me and Orson Welles released in theaters?
Answer: The film was released in the US on November 25, 2009.
U3
Model Question Relative position
Target What were critics reviews of Me and Orson Welles? 2 (U5)
FT When was the film Me and Orson Welles released in the UK? 0 (U3)
SOD What was the response to Me and Orson Welles? 2 (U5)
Table 14: An example of summarization on Newsroom. U1, U2, U7 and U8 are the 1st, 2nd, 7th and 8th utterances in the document. ‘Target’ is the target summary of the document and ‘FT’ and ‘SOD’ are the generated summary from FT and SOD. ‘Position’ is the position of the utterance associated with the summary in the document.
ID Document
U1 Joe Staley celebrates Colin Kaepernick’s touchdown run in Super Bowl XLVII.
U2 Cubs inspire Super Bowl memories for Staley Cubs fans were euphoric Wednesday night.
U7 Staley, 32, is the longest-tenured member of a team that’s 6-17 since 2015.
U8 Does he think the 49ers, who are nearly 22 years removed from their last title, can win a Super Bowl before he retires? “ Yeah,” he said, “I’ll never give up, man.”
Model Summary Position
Target Cubs inspire Super Bowl memories for Staley Cubs fans were euphoric Wednesday night U2
FT Joe Staley celebrates Colin Kaepernick’s touchdown run in Super Bowl XLVII. Cubs inspire Super Bowl memories for Staley. U1
SOD Joe Staley says he wished the 49ers had won the Super Bowl after the 2012 season. U8
Table 15: An example of summarization on CNN/DM. U1–U3 are the first three utterances in the document. ‘Target’ is the target summary of the document and ‘FT’ and ‘SOD’ are the generated summary from FT and SOD. ‘Position’ is the position of the utterance associated with the summary in the document.
ID Document
U1 Negotiations between the United States and Libya that could result in compensation for past acts of state-sponsored terrorism by Libya are under way.
U2 The wreckage of Pan Am 103 in Lockerbie, Scotland; the bombing killed 270 people in 1989. U.S. and Libyan officials met Wednesday and Thursday, the official said.
U3 The nations hope to hammer out a deal in which Libya would “resolve all outstanding claims in good faith” and offer “fair compensation” to victims and their families, he said.
Model Summary Position
Target The negotiations could result in compensation for past acts of state-sponsored terrorism by Libya. U3
FT The nations hope to hammer out a deal in which Libya would “resolve all outstanding claims in good faith”. U1
SOD Nations hoping for deal in which Libya would compensate terrorism victims. U3
Table 16: An example of KGC on Doc2dial. U7–U10 represent the consecutive four utterances following the 7th utterance in the document. T1 and T2 are the 1st and 2nd turn utterances of the dialogue. ‘Target’ is the target response of the example, ‘FT’ and ‘SOD’ are the generated responses from FT and SOD. ‘Position’ is the position of the grounded utterance of the response in the document and ‘Relative position’ is the distance of grounded utterances between the current response and the last turn response.
ID Document
U7 Should I Sign Up For Medical Insurance (Part B)?
U8 When you sign up for Medicare, you will be asked if you want to enroll in Medical Insurance Part B.
U9 If you do not choose to enroll in Medicare Part B and then decide to do so later, your coverage maybe delayed and you may have to pay a higher monthly premium unless you qualify for a “Special Enrollment Period” or SEP.
U10 You may qualify for a SEP that will let you sign up for Part B during …
ID Context Position
T1 Question: We don’t need to get any supplementary medical insurance, isn’t?
Answer: Yes, that’s correct. You may qualify for a SEP that will let you sign up for Part B.
U10
T2 Question: Do I sign up for medicare part B?
Model Answer Relative position
Target When you sign up for Medicare you will be asked if you want to enroll in Medical insurance Part B. -2 (U8)
FT Yes, that’s correct. You may qualify for a SEP that will let you sign up for Part B.    0 (U10)
SOD Yes, you will be asked if you want to enroll in Medical insurance Part B. -2 (U8)
Table 17: An example of KGC on Mutual. U1–U4 represent the four candidate responses. T1 and T2 are the 1st and 2nd turn utterances of the dialogue. ‘Target’ is the target response of the example, ‘FT’ and ‘SOD’ are the generated responses from FT and SOD. ‘Position’ is the position of the selected response.
ID Document
U1 It does n’t matter. you just joined a new team, and the manager said it’s normal that you are not good at interpersonal skills.
U2 Although the manager said you are not good at interpersonal skills, you still evaluated others’ performances.
U3 So you had your performance evaluation yesterday and were praised by the manager, right?
U4 yeah, you were praised by the manager, weren’t you?
ID Context Position
T1 Female: you look happy.
Male: I am. I had my performance evaluation today.
T2 Female: so it went well?
Male: yes, the manager said my interpersonal skills are great. I work well with others.
Model Female Position
Target Yeah, you were praised by the manager, weren’t you? U4
FT It does n’t matter. you just joined a new team, and the manager said it’s normal that you are not good at interpersonal skills. U1
SOD Yeah, you were praised by the manager, weren’t you? U4